# PRAGA: Prototype-aware Graph Adaptive Aggregation for Spatial Multi-modal Omics Analysis

### v3

@@ Abstract
Spatial multi-modal omics technology, highlighted by Nature Methods as an advanced biological technique in 2023, plays a critical role in resolving biological regulatory processes with spatial context. Recently, graph neural networks based on K-nearest neighbor (KNN) graphs have gained prominence in spatial multi-modal omics methods due to their ability to model semantic relations between sequencing spots. However, the fixed KNN graph fails to capture the latent semantic relations hidden by the inevitable data perturbations during the biological sequencing process, resulting in the loss of semantic information. In addition, the common lack of spot annotation and class number priors in practice further hinders the optimization of spatial multi-modal omics models. Here, we propose a novel spatial multi-modal omics resolved framework, termed PRototype-Aware Graph Adaptative Aggregation for Spatial Multi-modal Omics Analysis (PRAGA). PRAGA constructs a dynamic graph to capture latent semantic relations and comprehensively integrate spatial information and feature semantics. The learnable graph structure can also denoise perturbations by learning cross-modal knowledge. Moreover, a dynamic prototype contrastive learning is proposed based on the dynamic adaptability of Bayesian Gaussian Mixture Models to optimize the multi-modal omics representations for unknown biological priors. Quantitative and qualitative experiments on simulated and real datasets with 7 competing methods demonstrate the superior performance of PRAGA.
@@ Introduction
Spatially resolved transcriptomics, also refer as spatial transcriptomics, was crown as the Method of the Year by Nature (Xiaowei 2021). This technology expended the biological view of the gene expression abundance in single cells to spatial context, deciphering cell types and the heterogeneity in complex tissues. Recently, spatially resolved multi-modal omics, including transcriptomics, proteomics, and chromatin accessibility, was proposed to aggregate these modalities for comprehensively resolving gene regulation and microenvironment with spatial information in complex tissues (Li et al. 2024; Long et al. 2024).
The major challenge in spatial multi-modal omics is how to encode omics features from different modalities with corresponding spatial information into a unified latent space. Existing methods mainly build K-nearest neighbor (KNN) graphs to model the feature correlation with spatial positions between sequencing spots and generate unified comprehensive representations through Graph Neural Networks. For instance, spatial transcriptomic method STAGATE (Dong and Zhang 2022) utilizes KNN to construct a spatial adjacency graph for encoded transcriptomics data. Spatial multi-modal omics method SpatialGlue (Long et al. 2024) constructs the omics KNN graph as well as spatial adjacency graph for each modality separately and obtains a comprehensive latent representation through GCN (Kipf and Welling 2016). However, these methods ignore the interference in semantic relations caused by the perturbations inevitably introduced during the biological sequencing process, which KNN fails to overcome due to the artificially set K value limit.
@@
Multi-modal omics aggregation aims to integrate multiple omics data from the same biological sample to analyze gene expression and regulatory processes comprehensively. Existing multi-modal omics aggregation methods can be mainly divided into three categories: 1) non-negative matrix factorization; 2) Bayesian statistics; and 3) deep learning methods. The non-negative matrix factorization decomposes multi-modal omics data into a common factor matrix to obtain a unified representation (Jin, Zhang, and Nie 2020; Zeng et al. 2019). Bayesian statistics methods fit one omics data to the conditional probability distribution of other omics but require biological priors (Kamimoto et al. 2023; Bachireddy et al. 2021). Compared with the above two methods, deep learning methods have attracted widespread attention due to their scalability and excellent performance. TotalVI (Gayoso et al. 2021) jointly models RNA and protein data into a low-dimensional space based on a variational autoencoder to obtain the comprehensive representation. MultiVI (Ashuach et al. 2023) performs independent encoders for each modality of omics data and maps the multi-modal encodings to a joint latent space. Despite integrating multi-modal omics, these methods ignore the impact of spatial location on omics features.
@@
Recently, spatial resolved omics technologies, represented by spatial transcriptomics (Moses and Pachter 2022; Wei et al. 2022), have been proposed to associate spatial information with transcriptomic data. STAGATE (Dong and Zhang 2022) captures the local structure and spatial dependency of transcriptomic data via a graph attention network (Veličković et al. 2017). PAST (Li et al. 2023) performs a transformer architecture to capture self-similarity and global dependencies in transcriptomic data. The most advanced work SpatialGlue (Long et al. 2024) constructs a K-nearest neighbor (KNN) graph to model semantic relations and integrate spatial information with multi-modal omics through attention weights to provide a comprehensive representation. However, due to the inherent perturbations introduced during sequencing, the semantics of some omics features are disturbed and difficult to model by KNN graphs. In this paper, we propose a dynamic graph to mitigate the interference of perturbations and capture latent semantic relations through cross-modal knowledge and dynamic prototype contrastive learning.
@@
Equipped with the dynamic omic-specific graph, we construct a spatial aggregation graph to encode omics data by combining feature and spatial information. Taking RNA sequencing data $\boldsymbol{F_{\text{RNA}}}$ in transcriptomics as an example, we first initialize the learnable feature graph $\mathcal{G}^{F}_{\text{RNA}}$ and spatial adjacency graph $\mathcal{G}^{S}=(\boldsymbol{F_{\text{RNA}}},\boldsymbol{A^{S}})$ through KNN, where $\boldsymbol{A^{S}}\in\mathbb{R}^{N\times N}$ is the spatial adjacency matrix calculated based on spatial coordinates. The spatial aggregated graph $\hat{\mathcal{G}}_{\text{RNA}}=(\boldsymbol{F_{\text{RNA}}},\boldsymbol{\hat{A}_{\text{RNA}}})$ is obtained by combining $\mathcal{G}^{S}$ and $\mathcal{G}^{F}_{\text{RNA}}$ with learnable parameters:
$\boldsymbol{\hat{A}_{\text{RNA}}}=w^{S}_{\text{RNA}}\boldsymbol{A^{S}}+w^{F}_{\text{RNA}}\boldsymbol{A^{F}_{\text{RNA}}},$ (2)
where $w^{S}_{\text{RNA}}$ and $w^{F}_{\text{RNA}}$ are learnable parameters. Then, we perform a one-layer GCN (Kipf and Welling 2016) as an encoder to encode RNA sequencing features $\boldsymbol{F_{\text{RNA}}}$ with spatial aggregated graph $\hat{\mathcal{G}}_{\text{RNA}}$:
$\boldsymbol{Z_{\text{RNA}}}=\text{GCN}^{en}_{\text{RNA}}(\boldsymbol{F_{\text{RNA}}},\boldsymbol{\hat{A}_{\text{RNA}}})=\boldsymbol{\hat{A}_{\text{RNA}}}\boldsymbol{F_{\text{RNA}}}\boldsymbol{W_{\text{RNA}}^{en}},$ (3)
where $\boldsymbol{W_{\text{RNA}}^{en}}$ is a learnable parameter matrix in the encoder.
Similarly, we obtain the encoding of other omics data, e.g. $\boldsymbol{Z_{\text{ADT}}}$ and $\boldsymbol{Z_{\text{ATAC}}}$, using Eq. (2) and Eq. (3). After obtaining modality-specific encoding for all modalities $\{\boldsymbol{Z_{\text{RNA}}},\boldsymbol{Z_{\text{ADT}}},\boldsymbol{Z_{\text{ATAC}}}\}$, a Multi-Layer Perceptron (MLP) is utilized to map differential modality encoding to a unified comprehensive representation:
@@
[…]
@@
We conduct quantitative and qualitative experiments on five public datasets to verify the effectiveness of the proposed method: 1) Human Lymph Node dataset (Long et al. 2024); 2) Spatial epigenome–transcriptome mouse brain dataset (Zhang et al. 2023); 3) Mouse thymus stereo-CITE-seq dataset (Liao et al. 2023); 4) SPOTS mouse spleen dataset (Ben-Chetrit et al. 2023); 5) Spatial multi-modal omics simulation datasets (Long et al. 2024). These datasets are detailed in the Appendix.
@@
Table 3 shows the quantitative performance of PRAGA with different initial cluster numbers. Consistent with intuition, PRAGA performs best when the number of initial clusters is the same as the number of categories in Ground Truth. Nevertheless, when the number of initial clusters differs from Ground Truth, which is also a common case in practice, PRAGA still performs well without significant performance degradation. This demonstrates that PRAGA is reliable even if the category priors of sequencing points are unknown. Figure 4 shows the effect of different values of exponential moving average speed $\alpha$, temperature $\tau$ in $\mathcal{L}_{dpcl}$, and weight $\beta$ for $\mathcal{L}_{dpcl}$ on PRAGA performance. The experimental results consistently show that PRAGA is insensitive to the values of hyperparameters.
@@ Acknowledgements
The computational resources are supported by SongShan Lake HPC Center (SSL-HPC) in Great Bay University.

### v5

@@ Abstract
Spatial multi-modal omics technology, highlighted by Nature Methods as an advanced biological technique in 2023, plays a critical role in resolving biological regulatory processes with spatial context. Recently, graph neural networks based on K-nearest neighbor (KNN) graphs have gained prominence in spatial multi-modal omics methods due to their ability to model semantic relations between sequencing spots. However, the fixed KNN graph fails to capture the latent semantic relations hidden by the inevitable data perturbations during the biological sequencing process, resulting in the loss of semantic information. In addition, the common lack of spot annotation and class number priors in practice further hinders the optimization of spatial multi-modal omics models. Here, we propose a novel spatial multi-modal omics resolved framework, termed PRototype-Aware Graph Adaptative Aggregation (PRAGA). PRAGA constructs a dynamic graph to capture latent semantic relations and comprehensively integrate spatial information and feature semantics. The learnable graph structure can also denoise perturbations by learning cross-modal knowledge. Moreover, a dynamic prototype contrastive learning is proposed based on the dynamic adaptability of Bayesian Gaussian Mixture Models to optimize the multi-modal omics representations for unknown biological priors. Quantitative and qualitative experiments on simulated and real datasets with 7 competing methods demonstrate the superior performance of PRAGA. Code is available at https://github.com/Xubin-s-Lab/PRAGA.
@@ Introduction
Spatially resolved transcriptomics was crown as the Method of the Year by Nature (Xiaowei 2021). This technology expanded the biological view of the gene expression abundance in single cells to spatial context, deciphering cell types and the heterogeneity in complex tissues. Recently, spatially resolved multi-modal omics, including transcriptomics, proteomics, and chromatin accessibility, were proposed to aggregate these modalities for comprehensively resolving gene regulation and microenvironment with spatial information in complex tissues (Li et al. 2024; Long et al. 2024).
The major challenge in spatial multi-modal omics is how to encode omics features from different modalities with corresponding spatial information into a unified latent space. Existing methods mainly build K-nearest neighbor (KNN) graphs to model the feature correlation with spatial positions between sequencing spots and generate unified comprehensive representations through Graph Neural Networks. For instance, the spatial transcriptomic method STAGATE (Dong and Zhang 2022) utilizes KNN to construct a spatial adjacency graph for encoded transcriptomics data. Spatial multi-modal omics method SpatialGlue (Long et al. 2024) constructs the omics KNN graph as well as spatial adjacency graph for each modality separately and obtains a comprehensive latent representation through GCN (Kipf and Welling 2016). However, these methods ignore the interference in semantic relations caused by the perturbations inevitably introduced during the biological sequencing process, which KNN fails to overcome due to the artificially set K value limit.
@@
Multi-modal omics aggregation aims to integrate multiple omics data from the same biological sample to analyze gene expression and regulatory processes comprehensively (Wu et al. 2022; Zheng et al. 2024). Existing multi-modal omics aggregation methods can be mainly divided into three categories: 1) non-negative matrix factorization; 2) Bayesian statistics; and 3) deep learning methods. The non-negative matrix factorization decomposes multi-modal omics data into a common factor matrix to obtain a unified representation (Kim et al. 2020). Bayesian statistics methods fit one omics data to the conditional probability distribution of other omics but require biological priors (Argelaguet et al. 2020). Compared with the above two methods, deep learning methods have attracted widespread attention due to their scalability and excellent performance. TotalVI (Gayoso et al. 2021) jointly models RNA and protein data into a low-dimensional space based on a variational autoencoder to obtain the comprehensive representation. MultiVI (Ashuach et al. 2023) performs independent encoders for each modality of omics data and maps the multi-modal encodings to a joint latent space. Despite integrating multi-modal omics, these methods ignore the impact of spatial location on omics features.
@@
Recently, spatial resolved omics technologies, represented by spatial transcriptomics (Moses and Pachter 2022), have been proposed to associate spatial information with transcriptomic data. STAGATE (Dong and Zhang 2022) captures the local structure and spatial dependency of transcriptomic data via a graph attention network (Veličković et al. 2017). PAST (Li et al. 2023) performs a transformer architecture to capture self-similarity and global dependencies in transcriptomic data. The most advanced work SpatialGlue (Long et al. 2024) constructs a K-nearest neighbor (KNN) graph to model semantic relations and integrate spatial information with multi-modal omics through attention weights to provide a comprehensive representation. However, due to the inherent perturbations introduced during sequencing, the semantics of some omics features are disturbed and difficult to model by KNN graphs. In this paper, we propose a dynamic graph to mitigate the interference of perturbations and capture latent semantic relations through cross-modal knowledge and dynamic prototype contrastive learning.
@@
[…]
$\boldsymbol{\hat{A}_{m}}=w^{S}_{m}\boldsymbol{A^{S}}+w^{F}_{m}\boldsymbol{A^{F}_{m}},$ (2)
where $w^{S}_{m}$ and $w^{F}_{m}$ are learnable parameters. Then, we perform a one-layer GCN (Kipf and Welling 2016) as an encoder to encode sequencing features $\boldsymbol{F_{m}}$ with spatial aggregated graph $\hat{\mathcal{G}}_{m}$:
$\boldsymbol{Z_{m}}=\text{GCN}^{en}_{m}(\boldsymbol{F_{m}},\boldsymbol{\hat{A}_{m}})=\boldsymbol{\hat{A}_{m}}\boldsymbol{F_{m}}\boldsymbol{W_{m}^{en}},$ (3)
where $\boldsymbol{W_{m}^{en}}$ is a learnable parameter matrix in the encoder.
After obtaining modality-specific encoding for all modalities $\{\boldsymbol{Z_{\text{RNA}}},\boldsymbol{Z_{\text{ADT}}},\boldsymbol{Z_{\text{ATAC}}}\}$, a Multi-Layer Perceptron (MLP) is utilized to map differential modality encoding to a unified comprehensive representation:
@@
The reconstructed loss can be calculated by the mean square error between reconstructed omics features and origin omics features:
@@
We conduct quantitative and qualitative experiments on five public datasets to verify the effectiveness of the proposed method: 1) Human Lymph Node dataset (Long et al. 2024); 2) Spatial epigenome–transcriptome mouse brain dataset (Zhang et al. 2023); 3) Mouse thymus stereo-CITE-seq dataset (Liao et al. 2023); 4) SPOTS mouse spleen dataset (Ben-Chetrit et al. 2023); 5) Spatial multi-modal omics simulation datasets (Long et al. 2024). These datasets are detailed in the Appendix [^1] https://arxiv.org/abs/2409.12728.
@@
Table 3 shows the quantitative performance of PRAGA with different initial cluster numbers. Consistent with intuition, PRAGA performs best when the number of initial clusters is the same as the number of categories in Ground Truth. Nevertheless, when the number of initial clusters differs from Ground Truth, which is also a common case in practice, PRAGA still performs well without significant performance degradation. Figure 4 shows the effect of different values of exponential moving average speed $\alpha$, temperature $\tau$ in $\mathcal{L}_{dpcl}$, and weight $\beta$ for $\mathcal{L}_{dpcl}$ on PRAGA performance. Additional sensitivity experiments and the discussion of runtime can be found in the appendix. The experimental results consistently show that PRAGA is insensitive to the values of hyperparameters.
@@ Acknowledgements
The project is supported by the National Natural Science Foundation of China (Grant No. 32300554 and No. 62406056), and in part by the Guangdong Provincial Key Laboratory of Mathematical and Neural Dynamical Systems (Grant No.2024B1212010004). The computational resources are supported by Songshan Lake HPC Center (SSL-HPC) at Great Bay University.